# Many-Body Localization in the Age of Classical Computing

### v2

@@ VII.1.2 Assessing the breakdown of ergodicity
Tensor network investigation [604] of disordered XXZ spin chain revealed a slow decay of the density autocorrelation function $C(t)$ at time scales of the order of $10^{3}$ tunneling times deep in the MBL regime, which in ED studies at $L=20$ emerges at $W>W^{*}(L=20)\approx 4$. The decay, accurately fitted by a power-law with a small exponent, was found at $W=8$, as shown in Fig. 11(a), and at $W=10$. Notably, since the value of $C(t)$ changes by a few per milles throughout the examined time interval, the autocorrelation function can be fitted equally accurately with a logarithmic in time dependence, $C(t)\propto\mathrm{const}-\log(t)$. Nevertheless, in this section, we stick to the power-law fitting of $C(t)$, which describes the time dynamics in the disordered spin chains in the broad regime of disorder strengths and time scales. The exponent $\beta$ governing the decay of $C(t)$, extracted with a variety of computational approaches and across several literature references for disordered XXZ spin chain, decreases exponentially with disorder strength $W$, $\beta\propto e^{-\gamma W}$, in a wide interval of $W\in[1,10]$, as shown in Fig. 11(b). This result is another manifestation of the abrupt slowdown of the dynamics of disordered spin chains with disorder strength $W$. It is quantitatively consistent with the exponential in $W$ increase of the Thouless time (32) found in the ED studies.
@@
In contrast, Scenario 2. assumes that both $W^{T}(L)$ and $W^{*}(L)$ diverge, for instance proportionally to system size $L$, as suggested by [690, 572], in the $L\to\infty$ limit. In that case, no MBL phase exists in the considered disordered spin chains, and ETH (4) applies at any disorder strength $W$ for $L\to\infty$. Within this Scenario, the decay of $C(t)$ observed at times $t>10^{2}$ for $W=8$ in Fig. 11(a), and for $W=10$ in [604], could be a stable feature of the dynamics that persists until the time scales of the order of Heisenberg time, leading to thermalization of the system at sufficiently large system size $L$.
@@
Presently available numerical results do not allow us to distinguish the two Scenarios or point toward any other type of behavior. The decay of $C(t)$ at $W=8$ and $10$ could remain algebraic at longer times, leading to the thermalization of the system. However, we cannot exclude that $C(t)$ decays slower than the power-law on longer time scales and is only well fitted by a power-law in relatively narrow intervals of times $10^{2}<t<10^{3}$. Faced with an alternative of two competing ideas, we could ask which of the Scenarios is simpler given the data at our disposal. Examining the trends in the decay of $C(t)$ for system sizes $12\leq L\leq 20$ at times much smaller than the Heisenberg time $t_{H}$ and $W=3,5$, shown in Fig. 10(b), 10(c), respectively, we may conclude that the increase of system size always leads to a quicker decay that persists to a longer time. From that perspective, demanding that the decay of $C(t)$ presented in Fig. 11(a) slows down at a time larger than $t\approx 10^{3}$ but still much smaller than $t_{H}$, seems equivalent to requiring an emergence of a new trend in the data, hence placing an additional assumption on the dynamics of the system. On the other hand, the $C(t)$ decays only by a small fraction of its value in the available time interval $10^{2}<t<10^{3}$. Moreover, the shift of the crossing point $W^{*}(L)$ does slow down with $L$ for system sizes available to ED studies. Demanding that the slowdown of that drift ceases to occur at sufficiently large $L$ is yet another assumption about the system in a regime that we cannot presently probe.
@@
The above discussion illustrates that the attempts to answer the question of the existence of the MBL phase based solely on the presently available numerical data belong to the realm of speculation. The exponential slowdown of the dynamics of spin chains with increasing disorder strength is the root of the problems, preventing us from deciding whether the dynamics slow down but the thermalization process continues indefinitely or whether the dynamics get ultimately arrested at sufficiently large disorder strengths and system sizes. Irrespective of the answer to the question of the MBL phase, the numerically observed MBL regime is very robust in the sense that the time dynamics at large $W$ are extremely slow. For instance, if the decay of $C(t)$ continued to be an algebraic decay with the exponent $\beta\approx 10^{-3}$ observed for $W=8$, the $C(t)$ would decay to $10\%$ of its initial value after time $t\approx 10^{1000}$ tunneling times. In that case, the dynamics would effectively be arrested for all practical purposes.
@@ VII.1.3 Spatial decay of correlations
The density autocorrelation function $C(t)$ discussed above provides a scalar quantity that probes the approach to thermal equilibrium of the disordered spin chains. A connected two-point correlation function $\Phi(x,t)\equiv\langle S^{z}_{j+x}(t)S^{z}_{j}(0)\rangle-\langle S^{z}_{j+x}(t)\rangle\langle S^{z}_{j}(0)\rangle$ allows to probe the spatial correlations in the system, yielding much more complete information about the dynamics. A hydrodynamic description of diffusive dynamics [415] predicts that
@@
where $F_{0,0}$, $F_{1,0}$ are scaling functions of the scaling variable $y\equiv x/\sqrt{Dt}$ with diffusion constant $D$, and $d_{0}$ is a constant. The leading scaling function $F_{0,0}(y)=e^{-y^{2}/4}$ solves the diffusion equation, while the sub-leading terms include perturbative corrections [122] derived within the effective field theory of dissipative fluids [143]. For subdiffusive classical dynamics the scaling functions can present significant subleading corrections, as shown recently in [409]. Even in classical systems, where the system size can be pushed to the thousands of spins, the determination of the diffusion coefficient, or the subdiffusion exponent, must pass through the correct modellization of the subleading corrections.
The two-point correlation function $\Phi(x,t)$ was probed with the Chebyshev time propagation algorithm in [55] in the disordered XXZ spin chain of size $L\leq 24$. The analysis found a pronounced non-Gaussian shape of $\Phi(x,t)$, which decays nearly exponentially in $x$ even in the ergodic regime at $W=1.5<W^{T}(L)$ [702]. Analysis of the time dependence of the width of the $\Phi(x,t)$ profile yields a dynamical exponent $z$ significantly larger than the diffusive value $z=2$. However, at longer times, the value of the exponent $z$ becomes strongly dependent on the system size, decreasing towards $z=2$ with an increase of $L$. This prevents us from concluding whether the observed subdiffusive dynamics are a stable property of disordered XXZ spin chains or whether the features of subdiffusion are transient and disappear at sufficiently large times and system sizes, being replaced by diffusive dynamics. We revisit this question in Sec. VIII.
@@ VII.1.4 Overview
Time dynamics of local observables is a direct approach to assessing the ergodicity and class of dynamics in interacting many-body systems. Probing of time evolution is not only directly relevant for experiments with synthetic quantum matter but also numerically feasible, in limited time intervals, for system sizes larger than those accessible in ED studies.
@@ VII.2 Entanglement growth
The growth of entanglement in quench protocols provides another perspective on ergodicity and MBL in disordered many-body systems. While probing of entanglement entropy is more challenging than measurements of correlation functions of local observables, see, e.g., [191], entanglement entropy growth in disordered many-body system has been experimentally observed in [397]. In the following, we describe the properties of slow entanglement spreading in the MBL regime. We highlight the trends toward thermalizing behavior, which prevent us from pinpointing the status of the MBL phase, similarly as in the case of the autocorrelation function of local observables.
@@ VII.2.1 Logarithmic growth of entanglement entropy
Understanding the growth of entanglement entropy under unitary dynamics of many-body systems has been a subject of intense efforts over the last two decades. In non-interacting systems [102], the entanglement entropy $S(t)$, defined by Eq. (38), increases linearly with time $t$. Such a ”ballistic” growth of $S(t)$ is also found in generic ergodic systems [324]. The introduction of symmetries may alter this behavior. For instance, the Rényi entanglement entropy grows diffusively, proportionally to $\sqrt{t}$, in ergodic models with $U(1)$ symmetry [517] [^26] This behavior was substantiated in [283] and confirmed numerically in [264]., which, however, is specific to qubit systems in 1D [736]. General properties of the entanglement growth in higher dimensional systems, including the leading term quantifying the average value of $S(t)$ and sub-leading terms describing the entanglement fluctuations, are predicted by the minimal membrane picture conjectured in [437] and substantiated numerically beyond 1D in [599].
@@
Deep in the MBL regime of disordered XXZ spin chain, as exemplified in Fig. 12(a) for $W=10$, a logarithmic time dependence $S(t)\sim\log(t)$ accurately describes the entanglement entropy growth in the interval of times $t\in[200,4000]$. However, examining the entanglement entropy behavior in a broader time regime starting at $t\approx 10$, we see an evident curvature of the $S(t)$ on the log-linear scale. Consequently, a power-law increase $S(t)\sim t^{\gamma}$ with exponent $\gamma\ll 1$ describes the results rather well in such broad time interval. This behavior appears on the same time scales and system sizes as the slow decay of the density autocorrelation function $C(t)$ shown in Fig. 11(a). Hence, the better adherence of $S(t)$ to power-law fit is another manifestation of the tendency towards slow thermalization in the system. This conclusion must be, however, treated with caution. Firstly, we cannot exclude that the logarithmic time dependence could describe the increase of $S(t)$ better than the power-law at time scales longer than $t>4000$. Moreover, a comparison with predictions of the LIOM model of [131] reveals an apparent curvature of $S(t)$ on the log-linear scale in time interval $t\in[10,10^{4}]$. This behavior is entirely consistent with data presented in Fig. 11(a), while the LIOM model considered in [131] possesses, by construction, all features ascribed to the MBL phase, including the exponentially localized LIOMs and exponentially decaying interactions between them.
@@
Figure 12: Slow growth of entanglement in disordered XXZ spin chain. (a) Entanglement entropy $S(t)$ deep in the MBL regime, for $W=10$, calculated with TDVP for system size $L=50$, for bipartition of the chain into two contiguous subsystems each containing $L/2$ sites, averaged over more than $10^{3}$ disorder realizations. Power-law fit $S(t)\propto t^{\gamma}$ with $\gamma\approx 0.08$ described data accurately in a longer time interval than a logarithmic fit $S(t)\propto\log(t)$. (b) Number entropy $S_{n}(t)$ deep in the MBL regime is accurately captured by the double logarithmic fit $S(t)\propto\ln\ln(t)$. Reprinted and adapted from [604].
@@
A better understanding of the dynamics of LIOM models similar to the one considered by [131] will allow us to formulate more quantitative predictions on the form of the growth of the entanglement entropy $S(t)$ consistent with the MBL phase in the system. The asymptotic logarithmic growth, $S(t)\sim\log(t)$, of entanglement entropy, which is expected to be a signature of the MBL phase, is not easily distinguishable from an algebraic growth of $S(t)$, which would signify a thermalization of the system, again disallowing clear conclusions about the presence and the extent of the MBL phase. While we have concentrated here on entanglement entropy, the time evolution of different, more accessible entanglement measures, such as concurrence, has been investigated in the MBL regime in [290, 103]. Concurrence provides a spatiotemporal characterization of the entanglement dynamics in strongly disordered spin chains, revealing striking similarities with the dynamical heterogeneity in classical glasses [34].
@@ VII.2.2 Symmetry resolved entanglement: number entropy
The presence of globally conserved quantities allows a decomposition of the density matrix of a subsystem of the many-body system into orthogonal sectors, giving rise to the notion of symmetry resolved entanglement [235, 435, 668], and uncovering counter-intuitive behaviors such as the quantum Mpemba effect [32]. In particular, the conservation of the total $z$ spin component, $S^{z}\equiv\sum_{i=1}^{L}S^{z}_{i}$, i.e., the $U(1)$ symmetry of the disordered XXZ spin chain (7), opens up the possibility of studying a number entropy at the ETH-MBL crossover, providing another perspective on the dynamics of the disordered spin chains.
@@
We consider a bipartition of the 1D chain into subsystems $A$ and $B$, and denote $S^{z}_{A}=\sum_{i\in A}S^{z}_{i}$, $S^{z}_{B}=\sum_{i\in B}S^{z}_{i}$. The $U(1)$ symmetry implies that the total $z$ spin component commutes with the density matrix $\rho=\ket{\psi}\bra{\psi}$ | corresponding to the state of the system |$\ket{\psi}$, $[S^{z},\ket{\psi}\bra{\psi}]=0$. Taking the trace over the subsystem $B$ of the latter equation, and using the identities $\mathrm{tr}_{B}\left(S^{z}_{A}\rho\right)=\mathrm{tr}_{B}\left(\rho\right)S^{z}_{A}$ and $\mathrm{tr}_{B}\left(S^{z}_{B}\rho\right)=\mathrm{tr}_{B}\left(\rho S^{z}_{B}\right)$, we get
$[S^{z}_{A},\rho_{A}]=0,$ (49)
@@
implying that the reduced density matrix $\rho_{A}$ of the subsystem $A$ has a block diagonal structure, with blocks indexed by the eigenvalues $s\in[-L_{A}/2,\ldots,L_{A}/2]$ of the $S^{z}_{A}$ operator. Denoting by $\Pi_{s}$ a projector into the subspace of $S^{z}_{A}$ to eigenvalue $s$, we have
$\rho_{A}=\sum_{s}{p_{s}}\rho_{A,s},$ (50)
@@
where $p_{s}=\mathrm{tr}_{A}\left(\Pi_{s}\rho_{A}\Pi_{s}\right)$ and $\rho_{A,s}\equiv\Pi_{s}\rho_{A}\Pi_{s}/p_{A}$. The block structure in (50) allows us to rewrite the entanglement entropy (38) as $S=S_{n}+S_{c}$, where
$S_{n}=-\sum_{s}{p_{s}}\ln(p_{s}).$ (51)
@@
The quantity $S_{n}$ is known as the number entropy [562, 563, 176, 397], since it is related to the distribution of probabilities of finding the subsystem $A$ in various eigenspaces of the $S^{z}_{A}$ operator, which is, upon Jordan-Wigner transformation, equal to the particle number operator in the subsystem $A$ up to a constant shift. The quantity $S_{c}$, called configurational entanglement entropy, stems from superpositions of states from sector with fixed value of $S^{z}_{A}$ but with different spin configurations.
@@
For systems of free fermions, the number entropy $S_{n}$ can be used to bound the entanglement entropy $S$ from below and above, as derived in [323]. The bounds imply an asymptotic scaling $S\propto\exp\left(S_{n}\right)$, showing that growth of the entanglement entropy always implies a growth, although logarithmically slower, of the number entropy and vice versa [^27] The free fermion bounds in [323] were derived for Rényi entropy with index $q=2$, but may be expected to hold for any $q$.. These observations prompted an investigation of the number entropy in the strongly disordered XXZ spin chain [318].
@@
Numerical results of [318] for the disordered XXZ spin chain indicate, instead, a double logarithmic increase of the number entropy, $S_{n}(t)\propto\ln\ln(t)$. This behavior, consistent with the relation for non-interacting models $S\propto\exp\left(S_{n}\right)$, was confirmed by a TDVP study [604] for $L=50$, deep in the MBL regime of the disordered XXZ spin chain at $W=10$, see Fig. 12(b). The double logarithmic growth of $S_{n}(t)$ was interpreted [318] as a manifestation of persistent particle transport at a very slow rate, signifying the absence of the MBL phase. This leads to the proposition of an alternative picture of entanglement growth in MBL regime [319, 320], in which $S(t)\propto\ln(t)$ and $S_{n}(t)\propto\ln\ln(t)$.
@@
The observation of the double logarithmic growth of the number entropy prompted a study of $S_{n}(t)$ in a LIOM circuit model [131], which is, by construction, in MBL phase in the $L,t\to\infty$ limit. For parameters of the LIOM circuit that lead to the time dynamics parametrically similar to the disordered XXZ spin chain considered in [395], the LIOM circuit shows a clear growth $S_{n}(t)\propto\ln\ln(t)$. Notably, the conclusion about the double logarithmic growth of number entropy persists to the largest system size $L=28$ considered in [131], larger than the system size examined in [318]. The slow growth of $S_{n}(t)$ persists until a time that scales exponentially with $L$, similarly to the Heisenberg time $t_{H}$. This finding suggests that the ultraslow growth of the number entropy growth may be a feature inherent to the MBL phase [^28] The presently available results fo the LIOM circuits model do not fully exclude the possibility that the growth of $S_{n}(t)$ may saturate in LIOM circuits at larger times and system sizes.. Moreover, as suggested by [318], the $S_{n}(t)\propto\ln\ln(t)$ is reminiscent of the dynamics for which the time scale $t_{sp}$ for spreading of particles over length $\ell$ is exponentially large, $t_{sp}\propto e^{\ell}$. Depending on the factors in that relation, the time scale for reaching ergodicity, i.e., spreading particles over the entire system, $t_{sp}$ for $\ell=L$, may scale faster or slower than the Heisenberg time depending on the prefactors in the scaling. This may or may not enable a transition between ergodic and MBL phases in the disordered spin chains, and its resolution remains an open question.

### v3

@@ VII.1.2 Assessing the breakdown of ergodicity
Tensor network investigation [653] of the disordered XXZ spin chain revealed a slow decay of the density autocorrelation function $C(t)$ at time scales of the order of $10^{3}$ tunneling times deep in the MBL regime, which in ED studies at $L=20$ emerges at $W>W^{*}(L=20)\approx 4$. The decay, accurately fitted by a power-law with a small exponent, was found at $W=8$, as shown in Fig. 11(a), and at $W=10$. Notably, since the value of $C(t)$ changes by a few per milles throughout the examined time interval, the autocorrelation function can be fitted equally accurately with a logarithmic in time dependence, $C(t)\propto\mathrm{const}-\log(t)$. Nevertheless, in this Section, we stick to the power-law fitting of $C(t)$, which describes the dynamics in the disordered spin chains in the broad regime of disorder strengths and time scales. The exponent $\beta$ governing the decay of $C(t)$, extracted with a variety of computational approaches and across several literature references for the disordered XXZ spin chain, decreases exponentially with disorder strength $W$, $\beta\propto e^{-\gamma W}$, in a wide interval of $W\in[1,10]$, as shown in Fig. 11(b). This result is another manifestation of the abrupt slowdown of the dynamics of disordered spin chains with the disorder strength $W$. Arguments for classical subdiffusive systems relate the exponent $\beta$ with the dynamical exponent $z$ [425] via $\beta=1/z$. The exponential decay $\beta\propto e^{-\gamma W}$ is qualitatively consistent with the rapid increase of the exponent $z$ extracted from the analysis of Thouless time with the subdiffusive ansatz $t_{\mathrm{Th}}\propto L^{z}$ [748].
@@
In contrast, Scenario 2. assumes that both $W^{T}(L)$ and $W^{*}(L)$ diverge, for instance, proportionally to system size $L$, as suggested by [746, 619], in the $L\to\infty$ limit. In that case, no MBL phase exists in the considered disordered spin chains, and ETH (4) applies at any disorder strength $W$ for $L\to\infty$. Within this Scenario, the decay of $C(t)$ observed at times $t>10^{2}$ for $W=8$ in Fig. 11(a), and for $W=10$ in [653], could be a stable feature of the dynamics that persists until the time scales of the order of Heisenberg time, leading to thermalization of the system at sufficiently large system size $L$.
@@
Presently available numerical results do not allow us to distinguish the two Scenarios or point toward any other type of behavior. The decay of $C(t)$ at $W=8$ and $10$ could remain algebraic at longer times, leading to the thermalization of the system. However, we cannot exclude that $C(t)$ decays slower than the power-law on longer time scales and is only well fitted by a power-law in relatively narrow intervals of times $10^{2}<t<10^{3}$. Faced with an alternative of two competing ideas, we could ask which of the Scenarios is simpler given the numerical results at our disposal. Examining the trends in the decay of $C(t)$ for system sizes $12\leq L\leq 20$ at times much smaller than the Heisenberg time $t_{H}$ and $W=3,5$, shown in Fig. 10(b), 10(c), respectively, we may conclude that the increase of system size always leads to a quicker decay that persists to a longer time. From that perspective, demanding that the decay of $C(t)$ presented in Fig. 11(a) slows down at a time larger than $t\approx 10^{3}$ but still much smaller than $t_{H}$, seems equivalent to requiring an emergence of a new trend in the data, hence placing an additional assumption on the dynamics of the system. On the other hand, the $C(t)$ decays only by a small fraction of its value in the available time interval $10^{2}<t<10^{3}$. Moreover, the shift of the crossing point $W^{*}(L)$ does slow down with $L$ for system sizes available to ED studies. Demanding that the slowdown of that drift ceases to occur at sufficiently large $L$ is yet another assumption about the system in a regime that we cannot presently probe.
@@
The above discussion illustrates that the attempts to answer the question of the existence of the MBL phase based solely on the presently available numerical results belong to the realm of speculation. The exponential slowdown of the dynamics of spin chains with increasing disorder strength is the root of the problems, preventing us from deciding whether the dynamics slow down but the thermalization process continues indefinitely or whether the dynamics get ultimately arrested at sufficiently large disorder strengths and system sizes. Irrespective of the answer to the question of the MBL phase, the numerically observed MBL regime is very robust in the sense that the dynamics at large $W$ are extremely slow. For instance, if the decay of $C(t)$ continued to be an algebraic decay with the exponent $\beta\approx 10^{-3}$ observed for $W=8$, the $C(t)$ would decay to $10\%$ of its initial value after time $t\approx 10^{1000}$ tunneling times. In that case, the dynamics would effectively be arrested for all practical purposes.
@@ VII.1.3 Spatial decay of correlations
The density autocorrelation function $C(t)$ discussed above provides a scalar quantity that probes the approach to thermal equilibrium of the disordered spin chains. A connected two-point correlation function $\Phi(x,t)\equiv\langle\hat{S}^{z}_{j+x}(t)\hat{S}^{z}_{j}(0)\rangle-\langle\hat{S}^{z}_{j+x}(t)\rangle\langle\hat{S}^{z}_{j}(0)\rangle$ allows to probe the spatial correlations in the system, yielding much more complete information about the dynamics. A hydrodynamic description of diffusive dynamics [451] predicts that
@@
where $F_{0,0}$, $F_{1,0}$ are the scaling functions of the scaling variable $y\equiv x/\sqrt{Dt}$ with the diffusion constant $D$, and $d_{0}$ is a constant. The leading scaling function $F_{0,0}(y)=e^{-y^{2}/4}$ solves the diffusion equation, while the sub-leading terms include perturbative corrections [128], derived within the effective field theory of dissipative fluids [148]. For subdiffusive classical dynamics the scaling functions can present significant subleading corrections, as shown recently in [445]. Even in classical systems, where the system size can be pushed to the thousands of spins, the determination of the diffusion coefficient, or the subdiffusion exponent, must pass through the correct modellization of the subleading corrections.
The two-point correlation function $\Phi(x,t)$ was probed with the Chebyshev time propagation algorithm in [58] in the disordered XXZ spin chain of size $L\leq 24$. The analysis found a pronounced non-Gaussian shape of $\Phi(x,t)$, which decays nearly exponentially in $x$ even in the ergodic regime at $W=1.5<W^{T}(L)$ [758]. Analysis of the time dependence of the width of the $\Phi(x,t)$ profile yields a dynamical exponent $z$ significantly larger than the diffusive value $z=2$. However, at longer times, the value of the exponent $z$ becomes strongly dependent on the system size, decreasing towards $z=2$ with an increase of $L$. This prevents us from concluding whether the observed subdiffusive dynamics are a stable property of the disordered XXZ spin chains or whether the features of subdiffusion are transient and disappear at sufficiently large times and system sizes, being replaced by diffusive dynamics. We revisit this question in Sec. VIII.
@@ VII.1.4 Overview
The dynamics of local observables is a direct approach to assessing the ergodicity and the class of dynamics in interacting many-body systems. Probing of time evolution is not only directly relevant for experiments with synthetic quantum matter but also numerically feasible, in limited time intervals, for system sizes larger than those accessible in ED studies.
@@ VII.2 Entanglement growth
The growth of entanglement in quench protocols provides another perspective on ergodicity and MBL in disordered many-body systems. While probing the entanglement entropy is more challenging than measurements of correlation functions of local observables, see, e.g., [201], entanglement entropy growth in disordered many-body system has been experimentally observed in [433]. In the following, we describe the properties of slow entanglement spreading in the MBL regime. We highlight the trends toward thermalizing behavior, which prevent us from pinpointing the status of the MBL phase, similarly as in the case of the autocorrelation function of local observables.
@@ VII.2.1 Logarithmic growth of entanglement entropy
Understanding the growth of entanglement entropy under unitary dynamics of many-body systems has been a subject of intense efforts over the last two decades. In non-interacting systems [105], the entanglement entropy $S(t)$, defined by Eq. (38), increases linearly with time $t$. Such a ”ballistic” growth of $S(t)$ is also found in generic ergodic systems [350]. The introduction of symmetries may alter this behavior. For instance, the Rényi entanglement entropy grows diffusively, proportionally to $\sqrt{t}$, in ergodic models with $U(1)$ symmetry [558] [^36] This behavior was substantiated in [303] and confirmed numerically in [282]., which, however, is specific to qubit systems in one dimension [793]. General properties of the entanglement growth in higher dimensional systems, including the leading term quantifying the average value of $S(t)$ and sub-leading terms describing the entanglement fluctuations, are predicted by the minimal membrane picture conjectured in [475] and substantiated numerically beyond one dimension in [648].
@@
Deep in the MBL regime of the disordered XXZ spin chain, as exemplified in Fig. 12(a) for $W=10$, a logarithmic time dependence $S(t)\sim\log(t)$ accurately describes the entanglement entropy growth in the interval of times $t\in[200,4000]$. However, examining the entanglement entropy behavior in a broader time regime starting at $t\approx 10$, we see an evident curvature of the $S(t)$ on the log-linear scale. Consequently, a power-law increase $S(t)\sim t^{\gamma}$ with exponent $\gamma\ll 1$ describes the results rather well in such broad time interval. This behavior appears on the same time scales and system sizes as the slow decay of the density autocorrelation function $C(t)$ shown in Fig. 11(a). Hence, the better adherence of $S(t)$ to power-law fit is another manifestation of the tendency towards slow thermalization in the system. This conclusion must be, however, treated with caution. Firstly, we cannot exclude that the logarithmic time dependence could describe the increase of $S(t)$ better than the power-law at time scales longer than $t>4000$. Moreover, a comparison with predictions of the LIOM model of [136] reveals an apparent curvature of $S(t)$ on the log-linear scale in time interval $t\in[10,10^{4}]$. This behavior is entirely consistent with the results presented in Fig. 11(a), while the LIOM model considered in [136] possesses, by construction, all features ascribed to the MBL phase, including the exponentially localized LIOMs and exponentially decaying interactions between them.
@@
Figure 12: Slow growth of entanglement in the disordered XXZ spin-1/2 chain. (a) Entanglement entropy $S(t)$ deep in the MBL regime, for $W=10$, calculated with TDVP for system size $L=50$, for bipartition of the chain into two contiguous subsystems each containing $L/2$ sites, averaged over more than $10^{3}$ disorder realizations. Power-law fit $S(t)\propto t^{\gamma}$ with $\gamma\approx 0.08$ describes the results accurately in a longer time interval than a logarithmic fit $S(t)\propto\log(t)$. (b) Number entropy $S_{n}(t)$ deep in the MBL regime is accurately captured by the double logarithmic fit $S(t)\propto\ln\ln(t)$. Reprinted and adapted from [653].
@@
A better understanding of the dynamics of LIOM models similar to the one considered by [136] will allow us to formulate more quantitative predictions on the form of the growth of the entanglement entropy $S(t)$, consistent with the emergence of the MBL phase in the system. The asymptotic logarithmic growth, $S(t)\sim\log(t)$, of entanglement entropy, which is expected to be a signature of the MBL phase, is not easily distinguishable from an algebraic growth of $S(t)$, which would signify thermalization of the system, again disallowing clear conclusions about the presence and the extent of the MBL phase. While we have concentrated here on entanglement entropy, the time evolution of different, more accessible entanglement measures, such as concurrence, has been investigated in the MBL regime [314, 106]. Concurrence provides a spatiotemporal characterization of the entanglement dynamics in strongly disordered spin chains, revealing striking similarities with the dynamical heterogeneity in classical glasses [36].
@@ VII.2.2 Symmetry resolved entanglement: number entropy
The presence of globally conserved quantities allows a decomposition of the density matrix of a subsystem of the many-body system into orthogonal sectors, giving rise to the notion of symmetry resolved entanglement [251, 473, 721], and uncovering counter-intuitive behaviors such as the quantum Mpemba effect [34]. In particular, the conservation of the total spin $z$ component, $\hat{S}^{z}\equiv\sum_{i=1}^{L}\hat{S}^{z}_{i}$, i.e., the $U(1)$ symmetry of the disordered XXZ spin chain (7), opens up the possibility of studying a number entropy at the ETH-MBL crossover, providing another perspective on the dynamics of the disordered spin chains.
@@
We consider a bipartition of the chain into subsystems $A$ and $B$, and denote $\hat{S}^{z}_{A}=\sum_{i\in A}\hat{S}^{z}_{i}$, $\hat{S}^{z}_{B}=\sum_{i\in B}\hat{S}^{z}_{i}$. The $U(1)$ symmetry implies that the total spin $z$ component commutes with the density matrix $\rho=\ket{\psi}\bra{\psi}$ | corresponding to the state of the system |$\ket{\psi}$, $[S^{z},\ket{\psi}\bra{\psi}]=0$. Taking the trace over the subsystem $B$ of the latter equation, and using the identities $\mathrm{tr}_{B}\left(\hat{S}^{z}_{A}\rho\right)=\mathrm{tr}_{B}\left(\rho\right)\hat{S}^{z}_{A}$ and $\mathrm{tr}_{B}\left(\hat{S}^{z}_{B}\rho\right)=\mathrm{tr}_{B}\left(\rho\hat{S}^{z}_{B}\right)$, we get
$[\hat{S}^{z}_{A},\hat{\rho}_{A}]=0,$ (49)
@@
implying that the reduced density matrix $\hat{\rho}_{A}$ of the subsystem $A$ has a block diagonal structure, with blocks indexed by the eigenvalues $s\in[-L_{A}/2,\ldots,L_{A}/2]$ of the $\hat{S}^{z}_{A}$ operator. Denoting by $\Pi_{s}$ a projector into the subspace of $\hat{S}^{z}_{A}$ to eigenvalue $s$, we have
$\hat{\rho}_{A}=\sum_{s}{p_{s}}\hat{\rho}_{A,s},$ (50)
@@
where $p_{s}=\mathrm{tr}_{A}\left(\Pi_{s}\hat{\rho}_{A}\Pi_{s}\right)$ and $\hat{\rho}_{A,s}\equiv\Pi_{s}\hat{\rho}_{A}\Pi_{s}/p_{s}$. The block structure in (50) allows us to rewrite the entanglement entropy (38) as $S=S_{n}+S_{c}$, where
$S_{n}$ $=$ $-\sum_{s}{p_{s}}\ln(p_{s}),$
$S_{c}$ $=$ $-\sum_{s}{p_{s}}\mathrm{tr}_{A}\left(\hat{\rho}_{A,s}\ln(\hat{\rho}_{A,s})\right)\;.$ (51)
@@
The quantity $S_{n}$ is known as the number entropy [609, 610, 186, 433], since it is related to the distribution of probabilities of finding the subsystem $A$ in various eigenspaces of the $\hat{S}^{z}_{A}$ operator, which is, upon Jordan-Wigner transformation, equal to the particle number operator in the subsystem $A$, up to a constant shift. The quantity $S_{c}$, called configurational entanglement entropy, stems from superpositions of states from sector with fixed value of $\hat{S}^{z}_{A}$ but with different spin configurations.
@@
For systems of free fermions, the number entropy $S_{n}$ can be used to bound the entanglement entropy $S$ from below and above, as derived in [349]. The bounds imply an asymptotic scaling $S\propto\exp\left(S_{n}\right)$, showing that the growth of the entanglement entropy always implies a growth, although logarithmically slower, of the number entropy, and vice versa [^37] The free fermion bounds in [349] were derived for Rényi entropy with index $q=2$, but may be expected to hold for any $q$.. These observations prompted an investigation of the number entropy in the strongly disordered XXZ spin chain [344].
@@
Numerical results of [344] for the disordered XXZ spin chain indicate, instead, a double logarithmic increase of the number entropy, $S_{n}(t)\propto\ln\ln(t)$. This behavior, consistent with the relation for non-interacting models, $S\propto\exp\left(S_{n}\right)$, was confirmed by a TDVP study [653] for $L=50$, deep in the MBL regime of the disordered XXZ spin chain at $W=10$, see Fig. 12(b). The double logarithmic growth of $S_{n}(t)$ was interpreted [344] as a manifestation of persistent particle transport at a very slow rate, signifying the absence of the MBL phase. This leads to the proposition of an alternative picture of entanglement growth in MBL regime [345, 346], in which $S(t)\propto\ln(t)$ and $S_{n}(t)\propto\ln\ln(t)$.
@@
The observation of the double logarithmic growth of the number entropy prompted a study of $S_{n}(t)$ in a LIOM circuit model [136], which is, by construction, in MBL phase in the $L,t\to\infty$ limit. For parameters of the LIOM circuit that lead to the time dynamics parametrically similar to the disordered XXZ spin chain considered in [426], the LIOM circuit shows a clear growth $S_{n}(t)\propto\ln\ln(t)$. Notably, the conclusion about the double logarithmic growth of number entropy persists to the largest system size $L=28$ considered in [136], larger than the system size examined in [344]. The slow growth of $S_{n}(t)$ persists until a time that scales exponentially with $L$, similarly to the Heisenberg time $t_{H}$. This finding suggests that the ultraslow growth of the number entropy growth may be a feature inherent to the MBL phase, even though the presently available results for the LIOM circuits model do not fully exclude the possibility that the growth of $S_{n}(t)$ may saturate in LIOM circuits at larger times and system sizes. Moreover, as suggested by [344], the $S_{n}(t)\propto\ln\ln(t)$ is reminiscent of the dynamics for which the time scale $t_{sp}$ for spreading of particles over length $\ell$ is exponentially large, $t_{sp}\propto e^{\ell}$. Depending on the factors in that relation, the time scale for reaching ergodicity, i.e., spreading particles over the entire system, $t_{sp}$ for $\ell=L$, may scale faster or slower than the Heisenberg time depending on the prefactors in the scaling. This may or may not enable a transition between ergodic and MBL phases in the disordered spin chains, and its resolution remains an open question.